Write the slope-intercept equation for the line that passes through (5, 4) and (7, 8).
Answer by rfer(16322)

slope m=(8-4)/(7-5)
m=4/2 =2
-----------
4=2*5+b
4=10+b
b=-6
-----------
y=2x-6